OpenStudy (lgbasallote):

LGBARIDDLE How many unique integer solutions exist for a + b + c = 5 such that (a, b and c) \(\ge\) 0

OpenStudy (agent47):

an infinite number

OpenStudy (lgbasallote):

no. it's finite

OpenStudy (agent47):

example: 1.9+1.1+2 1.99+1.01+2

OpenStudy (agent47):

1.999+1.001+2

OpenStudy (agent47):

these are just the ones im getting with 1.9.., 1.0...1, and 2

OpenStudy (agent47):

unless you specify that a, b and c are integers

OpenStudy (lgbasallote):

uhh yeah they are integers

OpenStudy (agent47):

SEEE, ok now my solution doesn't work :/

OpenStudy (agent47):

a+b=5-c c has to be less than or equal to 5, since a+b cannot be negative.. so now list all the choices: c=0, a+b=5 c=1, a+b=4 c=2, a+b=3 c=3, a+b=2 c=4, a+b=1 c=5, a+b=0... now expanding on this...

OpenStudy (agent47):

c=0: a=0, b=5 a=1, b=4 a=2, b=3 a=3, b=2 a=4, b=1 a=5, b=0

OpenStudy (agent47):

c=1: a=0, b=4 a=1, b=3 a=2, b=2 a=3, b=1 a=4, b=0

OpenStudy (lgbasallote):

wow brute force

OpenStudy (agent47):

6+5+4+3+2+1

OpenStudy (agent47):

11+7+3=11+10=21?

OpenStudy (agent47):

Brute force is my fave method.. I am a gorilla after all!

OpenStudy (anonymous):

21 sets of numbers for a, b, and c. If a=0, then 6 sets for b and c (0,5 and 1,4 and 2,3 and 3,2 and 4,1 and 5,0). If a=1, then 5 sets for b and c (0,4 and 1,3 and 2,2 and 3,1 and 4,0), etc., until you get to a=5, and then there is only the set of 0,0 for b and c. Adding 6 + 5 + 4 + 3 + 2 + 1 = 21.